OpenStudy (marcon):

Which polynomial identity will prove that 65 = 81 - 16? Binomial Multiplication Difference of Cubes Sum of Cubes Difference of Squares

OpenStudy (campbell_st):

well what type of numbers are 81 and 16..?

OpenStudy (marcon):

I think they're whole or real number. they're not prime

OpenStudy (campbell_st):

ok.... these are both special numbers.... not primes.... a special type of composites... and look at the answer choices for clues...

OpenStudy (marcon):

I think either A or B....

OpenStudy (marcon):

I'm not sure what it is. @campbell_st

OpenStudy (campbell_st):

ok... cube are 1^3 = 1 2^3 = 8 3^3 = 27.... do you think those numbers of 81 and 16 will be in that pattern..?

OpenStudy (marcon):

nope

OpenStudy (campbell_st):

ok... so I'll eliminate 1... and you have eliminated 2... so we know 81 and 16 aren't cubes... any thoughts.... and the operation subtraction - can be called difference...

OpenStudy (marcon):

I would say "differences of squares" as my answer because 4^2 = 16 and 9^2 is 81. Is that correct?

OpenStudy (campbell_st):

that's correct.

OpenStudy (marcon):

Alright thanks for your help

OpenStudy (marcon):

Which polynomial identity will prove that 64 + 27 = 91? Difference of Squares Difference of Cubes Sum of Cubes Square of a binomial So would this one be "sum of cubes" ? @campbell_st